# Harmonizing SO(3)-Equivariance with Neural Expressiveness: a Hybrid Deep Learning Framework Oriented to the Prediction of Electronic Structure Hamiltonian

### v8

@@ Abstract
Deep learning for predicting the electronic structure Hamiltonian of quantum systems necessitates satisfying the covariance laws, among which achieving SO(3)-equivariance without sacrificing the non-linear expressive capability of networks remains unsolved. To navigate the harmonization between equivariance and expressiveness, we propose a deep learning method synergizing two distinct categories of neural mechanisms as a two-stage cascaded regression framework. The first stage corresponds to group theory-based neural mechanisms with inherent SO(3)-equivariant properties prior to the parameter learning process, while the second stage is characterized by a non-linear 3D graph Transformer network we propose featuring high capability on non-linear expressiveness. The novel combination lies in the point that, the first stage predicts baseline Hamiltonians with abundant SO(3)-equivariant features extracted, assisting the second stage in empirical learning of equivariance; and in turn, the second stage refines the first stage’s output as a fine-grained prediction of Hamiltonians using powerful non-linear neural mappings, compensating for the intrinsic weakness on non-linear expressiveness capability of mechanisms in the first stage. Our method enables precise, generalizable predictions while maintaining robust SO(3)-equivariance under rotational transformations, and achieves state-of-the-art performance in Hamiltonian prediction on six benchmark databases.
@@ 1 Introduction
However, applying deep learning techniques to the Hamiltonian prediction task continues to present substantial challenges, highlighting the evolving demands on deep learning technology. An ultra-high accuracy to $10^{-1}$ meV scale ($10^{-4}$ eV) is required to precisely calculate the down-stream physical quantities such as band structures. Furthermore, the fidelity of Hamiltonian predictions should not be confined to a specific coordinate system; rather, the results must demonstrate robust symmetry and generalizability across various choices of reference frames. This necessitates that deep learning methods capture the intrinsic symmetries of the Hamiltonian with respect to transformations of the coordinate system, ensuring the consistency with physical principles covariant to these transformations. In the context of Hamiltonian prediction, the key covariance principles are 3D translational invariance and rotational equivariance. By utilizing relative coordinates, 3D translational invariance has been effortlessly achieved, while 3D rotational equivariance, i.e., equivariance to the SO(3) group, remains a challenging target to guarantee. This difficulty arises because the Hamiltonian of each pair of atoms is usually high-dimensional, and its variation space under rotational disturbance is large. Consequently, it is difficult to cover the vast variability space they inhabit merely depending on parameter learning from discrete training samples. To address this, several works [27, 6] applied group theory-guaranteed feature descriptors and tensor operators assuring inherent SO(3)-equivariance prior to the data-driven parameter learning process. However, to guarantee such SO(3)-equivariance independent to specific network parameters, these methods did not permit the use of non-linear activation layers for SO(3)-equivariant features, i.e. those with degrees greater than 00. Given that the expressiveness capability of modern deep networks heavily relies on the extensive use of non-linear mappings, this restriction resulted in a significant drawback on network’s expressiveness capabilities. This, in turn, led to bottlenecks in fitting complex mappings, limiting the accuracy achievable in predicting Hamiltonians. This dilemma, is also broadly prevalent in other 3D machine learning tasks where SO(3)-equivariance is highlighted, as analyzed by Zitnick et al. [28].
@@
To harmonize SO(3)-equivariance and expressiveness for the prediction of electronic structure Hamiltonians, this paper proposes a novel two-stage regression framework that combines mechanisms boasting parameter-indep-endent, prior SO(3)-equivariance with mechanisms featuring flexibility in non-linear expressiveness, overcoming the respective challenges of each categories of mechanisms, i.e., the limited non-linear expressive capability for the former as well as the difficulty of learning SO(3)-equivariance from data for the latter, through effective complementary strategies. Specifically, the first regression stage corresponds to the neural mechanisms constructed based on group theory with prior equivariant properties of 3D atomic systems, predicting an approximate value of the Hamiltonian, with abundant SO(3)-equivariant features provided. In the second stage, a highly expressive graph Transformer network we design, with no restrictions on non-linear activations, takes over. This network dynamically learns the 3D structural patterns of the atomic systems, compensates for the expressiveness shortcomings of the first-stage network arising from limited non-linear mappings, and refines the Hamiltonian values predicted in the first stage to enhance accuracy. Although this stage might not possess a parameter-independent prior SO(3)-equivariance due to its non-linearity, it is capable of capturing SO(3)-equivariance through learning effective network parameters via the help of three pivotal mechanisms. First, instead of directly regressing the entire Hamiltonians, the second stage aims to refine the Hamiltonian predictions from the first stage with corrective adjustments in a cascaded manner. The scopes of adjustments are smaller, lowing down the difficulties on non-linear learning of SO(3)-equivariance. Second, the second-stage network incorporates covariant features, including SO(3)-equivariant features extracted by the first-stage network and SO(3)-invariant features engineered by geometric knowledge, with its inputs to assist in the implicit learning of SO(3)-equivariance. Third, as the core of Transformer, the attention mechanism has potentials to adapt to geometric condition variations such as coordinate transformations, through its dynamic weighting strategy. Collectively, the combination of the two categories of neural mechanisms in the two stages allows the framework to overcome the challenges of each individual mechanism and make precise, generalizable, and SO(3)-equivariant predictions, being much more effective than simply increasing the parameter count for the network from one stage and fine-tuning it alone.
@@ 2 Related Work
In the field of researches on equivariance to SO(3), approaches like DeepH [13] explored equivariance via a local coordinate strategy, which made inference within the fixed local coordinate systems built with neighboring atoms, then transferred the output according to equivariance rules to the corresponding global coordinates. However, due to a lack of in-depth exploration of SO(3)-equivariance at the neural mechanism level, this method faced challenges when the local coordinate system underwent rotational disturbances from non-rigid deformation, e.g. the inter-layer twist of bilayer structures. In contrast, some methods [20, 4, 5] considered SO(3)-equivariance from the perspective of intrinsic mechanisms of neural networks, developed group theory-informed equivariant operations, such as linear scaling, element-wise sum, direct sum, direct product, as well as the Clebsch-Gordan decomposition, effectively applying in tasks like molecular dynamics simulation [16, 14] and Hamiltonian prediction [19, 21, 6], where DeepHE3 [6] was validated as a SOTA method across diverse atomic systems. However, a common challenge across these methods lies in the fact that, to achieve inherent equivariance prior to the parameter learning process, they forbade the use of non-linear mappings for SO(3)-equivariant features, significantly limiting the network’s expressive potential and creating a bottleneck in generalization performance. Although these approaches tried to enhance expressiveness via a gated activation function, where SO(3)-invariant features undergone through non-linear activation layers were used as gating coefficients that were multiplied with SO(3)-equivariant features, this mechanism, viewed from the perspective of equivariant features, amounted to a linear operation and did not fundamentally improve their expressive capability. For these methods, this equivariance-expressiveness dilemma remains an unsolved problem.
@@ 4 Method
As shown in Fig. 1, to harmonize SO(3)-equivariance and expressiveness for the deep learning paradigm, we propose a novel framework effectively hybridizing networks in two regression stages, from which the first-stage network, a group theory-informed network possessing SO(3)-equivariance prior to the learning process, provides essential foundations to the second stage in mastering SO(3)-equivariance, whereas the second-stage network, with highly expressive non-linear mappings, enriches the expressiveness capabilities of the whole framework. The combination of these two stages not only enhances expressiveness but also ensures robust equivariance to rotations of reference frames, bringing accurate predictions for electronic structure Hamiltonians despite rotational transformations.
@@ 6 Time Complexity
7 The Distinction Between Our Method and Other Transformer Architectures with Efforts to SO(3)-equivariance
In the realm of achieving SO(3)-equivariance in transformer architectures, some efforts have been made like SE(3)-transformers [4], Equiformer [14], and Equiformerv2 [15]. These methods aimed to merge the robust expressiveness with SO(3)-equivariance within a single Transformer architecture and faced substantial challenges. SE(3)-transformers and Equiformer, for instance, resorted to the gated mechanism as the activation function of the Transformer to maintain equivariance, compromising the non-linear expressiveness for features with degree $l\geq 1$, a limitation thoroughly analyzed in Section 2 of our work. Equiformerv2, striving for enhanced expressiveness, decomposed features into weighted sums of spherical harmonic basis functions corresponding to discrete points on a unit sphere, allowing for non-linear operations on the discrete coefficients. However, on improving non-linear expressiveness, it shifted from continuous to discrete rotational equivariance, diluting the model’s ability to generalize across continuous rotations in the SO(3) group.
In contrast to struggling on achieving both SO(3)-equivariance and powerful non-linear expressiveness with a single Transformer architecture, we explore a complementary combination of neural mechanisms boasting parameter-independent prior SO(3)-equivariance with mechanisms featuring enhanced non-linear expressiveness, in a hybrid two-stage regression framework. From which, our Transformer architecture is designed to fully release non-linear expressiveness capability while capturing SO(3)-equivariance in the parameter learning process within the hybrid cascaded regression context. Experiments in 5.2 demonstrates that this hybrid approach is much more effective on harmonizing SO(3)-equivariance and expressiveness than merely increasing the parameter counts of a single neural mechanism, thus achieving SOTA performance. Moreover, our framework offers flexibility, allowing for the integration of other mechanisms with prior equivariance or great non-linear expressiveness capability in future works, enhancing the universality and scalability for deep learning solutions on regressing electronic structure Hamiltonians.

### v9

@@ Abstract
Deep learning for predicting the electronic structure Hamiltonian of quantum systems necessitates satisfying the covariance laws, among which achieving SO(3)-equivariance without sacrificing the non-linear expressive capability of networks remains unsolved. To navigate the harmonization between equivariance and expressiveness, we propose a deep learning method, namely HarmoSE, synergizing two distinct categories of neural mechanisms as a two-stage cascaded regression framework. The first stage corresponds to group theory-based neural mechanisms with inherent SO(3)-equivariant properties prior to the parameter learning process, while the second stage is characterized by a non-linear 3D graph Transformer network we propose featuring high capability on non-linear expressiveness. The novel combination lies in the point that, the first stage predicts baseline Hamiltonians with abundant SO(3)-equivariant features extracted, assisting the second stage in empirical learning of equivariance; and in turn, the second stage refines the first stage’s output as a fine-grained prediction of Hamiltonians using powerful non-linear neural mappings, compensating for the intrinsic weakness on non-linear expressiveness capability of mechanisms in the first stage. Our method enables precise, generalizable predictions while maintaining robust SO(3)-equivariance under rotational transformations, and achieves state-of-the-art performance in Hamiltonian prediction on six benchmark databases.
@@ 1 Introduction
However, applying deep learning techniques to the Hamiltonian prediction task continues to present substantial challenges, highlighting the evolving demands on deep learning technology. An ultra-high accuracy to $10^{-1}$ meV scale ($10^{-4}$ eV) is required to precisely calculate the down-stream physical quantities such as band structures. Furthermore, the fidelity of Hamiltonian predictions should not be confined to a specific coordinate system; rather, the results must demonstrate robust symmetry and generalizability across various choices of reference frames. This necessitates that deep learning methods capture the intrinsic symmetries of the Hamiltonian with respect to transformations of the coordinate system, ensuring the consistency with physical principles covariant to these transformations. In the context of Hamiltonian prediction, the key covariance principles are 3D translational invariance and rotational equivariance. By utilizing relative coordinates, 3D translational invariance has been effortlessly achieved, while 3D rotational equivariance, i.e., equivariance to the SO(3) group, remains a challenging target to guarantee. This difficulty arises because the Hamiltonian of each pair of atoms is usually high-dimensional, and its variation space under rotational disturbance is large. Consequently, it is difficult to cover the vast variability space they inhabit merely depending on parameter learning from discrete training samples. To address this, several works [27, 6] applied group theory-guaranteed feature descriptors and tensor operators assuring inherent SO(3)-equivariance prior to the data-driven parameter learning process. However, to guarantee such SO(3)-equivariance independent to specific network parameters, these methods highly restricted the use of non-linear activation layers for SO(3)-equivariant features, i.e. those with degrees greater than 00. Given that the expressiveness capability of modern deep networks heavily relies on the extensive use of non-linear mappings, this restriction resulted in a significant drawback on network’s expressiveness capabilities. This, in turn, led to bottlenecks in fitting complex mappings, limiting the accuracy achievable in predicting Hamiltonians. This dilemma, is also broadly prevalent in other 3D machine learning tasks where SO(3)-equivariance is highlighted, as analyzed by Zitnick et al. [28].
@@
To harmonize SO(3)-equivariance and expressiveness for the prediction of electronic structure Hamiltonians, this paper proposes a novel two-stage regression framework, namely HarmoSE, which combines mechanisms boasting parameter-indep-endent, prior SO(3)-equivariance with mechanisms featuring flexibility in non-linear expressiveness, overcoming the respective challenges of each categories of mechanisms, i.e., the limited non-linear expressive capability for the former as well as the difficulty of learning SO(3)-equivariance from data for the latter, through effective complementary strategies. Specifically, the first regression stage corresponds to the neural mechanisms constructed based on group theory with prior equivariant properties of 3D atomic systems, predicting an approximate value of the Hamiltonian, with abundant SO(3)-equivariant features provided. In the second stage, a highly expressive graph Transformer network we design, with no restrictions on non-linear activations, takes over. This network dynamically learns the 3D structural patterns of the atomic systems, compensates for the expressiveness shortcomings of the first-stage network arising from limited non-linear mappings, and refines the Hamiltonian values predicted in the first stage to enhance accuracy. Although this stage might not possess a parameter-independent prior SO(3)-equivariance due to its non-linearity, it is capable of capturing SO(3)-equivariance through learning effective network parameters via the help of three pivotal mechanisms. First, instead of directly regressing the entire Hamiltonians, the second stage aims to refine the Hamiltonian predictions from the first stage with corrective adjustments in a cascaded manner. The scopes of adjustments are smaller, lowing down the difficulties on non-linear learning of SO(3)-equivariance. Second, the second-stage network incorporates covariant features, including SO(3)-equivariant features extracted by the first-stage network and SO(3)-invariant features engineered by geometric knowledge, with its inputs to assist in the implicit learning of SO(3)-equivariance. Third, as the core of Transformer, the attention mechanism has potentials to adapt to geometric condition variations such as coordinate transformations, through its dynamic weighting strategy. Collectively, the combination of the two categories of neural mechanisms in the two stages allows the framework to overcome the challenges of each individual mechanism and make precise, generalizable, and SO(3)-equivariant predictions, being much more effective than simply increasing the parameter count for the network from one stage and fine-tuning it alone.
@@ 2 Related Work
In the field of researches on equivariance to SO(3), approaches like DeepH [13] explored equivariance via a local coordinate strategy, which made inference within the fixed local coordinate systems built with neighboring atoms, then transferred the output according to equivariance rules to the corresponding global coordinates. However, due to a lack of in-depth exploration of SO(3)-equivariance at the neural mechanism level, this method faced challenges when the local coordinate system underwent rotational disturbances from non-rigid deformation, e.g. the inter-layer twist of bilayer structures. In contrast, some methods [20, 4, 5] considered SO(3)-equivariance from the perspective of intrinsic mechanisms of neural networks, developed group theory-informed equivariant operations, such as linear scaling, element-wise sum, direct sum, direct product, as well as the Clebsch-Gordan decomposition, effectively applying in tasks like molecular dynamics simulation [16, 14] and Hamiltonian prediction [19, 21, 6], where DeepHE3 [6] was validated as a SOTA method across diverse atomic systems. However, a common challenge across these methods lies in the fact that, to achieve inherent equivariance prior to the parameter learning process, they forbade the use of complex non-linear mappings like $Sigmoid$, $SiLU$, and $Softmax$ for SO(3)-equivariant features, significantly limiting the network’s expressive potential and creating a bottleneck in generalization performance. Although these approaches tried to enhance expressiveness via a gated activation function, where SO(3)-invariant features undergone through non-linear activation layers were used as gating coefficients that were multiplied with SO(3)-equivariant features, this mechanism, viewed from the perspective of equivariant features, amounted to a linear operation and did not fundamentally improve their expressive capability. For these methods, this equivariance-expressiveness dilemma remains an unsolved problem.
@@ 4 Method
As shown in Fig. 1, to harmonize SO(3)-equivariance and expressiveness for the deep learning paradigm, we propose a novel framework, namely HarmoSE, effectively hybridizing networks in two regression stages, from which the first-stage network, a group theory-informed network possessing SO(3)-equivariance prior to the learning process, provides essential foundations to the second stage in mastering SO(3)-equivariance, whereas the second-stage network, with highly expressive non-linear mappings, enriches the expressiveness capabilities of the whole framework. The combination of these two stages not only enhances expressiveness but also ensures robust equivariance to rotations of reference frames, bringing accurate predictions for electronic structure Hamiltonians despite rotational transformations.